# Disentangling the galactic and intergalactic components in 313 observed Lyman-alpha line profiles between redshift 0 and 5

### v1

@@ Abstract
Context. Lyman-Alpha (${\rm Ly}\alpha$) photons emitted in star-forming regions inside galaxies experience a complex radiative transfer process until they reach the observer. The ${\rm Ly}\alpha$ line profile that we measured on Earth is, thus, the convolution of the gas properties in the interstellar (ISM), circumgalactic (CGM), and intergalactic medium (IGM).
Aims. We make use of the open source package zELDA (redshift Estimator for Line profiles of Distant Lyman-Alpha emitters) to disentangle the galactic and IGM components of the ${\rm Ly}\alpha$ profiles to study both the evolution of the intrinsic galactic emission and the IGM transmission across cosmic time.
Methods. zELDA includes different artificial neural networks that reconstruct IGM attenuated ${\rm Ly}\alpha$ line profiles. These models are trained using mock ${\rm Ly}\alpha$ line profiles. A Monte Carlo radiative transfer code computes the galactic component for the so-called thin shell model. Moreover, the IGM component is included through the IGM transmission curves generated from the IllustrisTNG100 cosmological galaxy formation simulation.
Results. We recover their intrinsic galactic spectra by applying the zELDA to 313 ${\rm Ly}\alpha$ line profiles observed with HST/COS and MUSE. Sources at z ¡ 0.5 show weak IGM attenuation, while at z ¿ 3, ZELDA reveals significant IGM suppression of the blue peak in several sources. After separating the IGM effects, the stacked intrinsic galactic ${\rm Ly}\alpha$ profiles show a minimal evolution from z = 0 to 6. The mean IGM transmission for z ¡ 0.5 in HST/COS data exceeds 90%, while the MUSE data show an evolution from 0.85 at z = 3.0 to 0.55 at z = 5.0.
Key Words.:
Radiative transfer, Galaxies: intergalactic medium, Galaxies: ISM
@@ 1 Introduction
Figure 1: Illustration of the impact of the IGM attenuation in the galactic ${\rm Ly}\alpha$ line profile. Through all the panels, we show the intrinsic ${\rm Ly}\alpha$ line profile emerging from the galaxy in thin black and the IGM transmission curve in yellow. In the top panels, we show the galactic line profile obscured (purple). The central left panels show the galactic line profile after traveling a short distance in the IGM (blue). As the ${\rm Ly}\alpha$ travel further through the IGM the line is more and more absorbed in the blue side (top right central panel), until they reach the observer (red). The bottom panels show the zELDA reconstruction (purple) using as input the observed line profile (red), along with the line structure terminology used in this work.
@@
Specifically, the shape of the observed ${\rm Ly}\alpha$ line profile is the product of the complex radiative transfer processes that take place in the interstellar, circumgalactic, and intergalactic mediums (ISM, CGM, and IGM, respectively; for a review on the radiative transfer see Dijkstra, 2017). The high cross-section between neutral hydrogen and ${\rm Ly}\alpha$ causes these photons to be absorbed and reemitted, often thousands of times before leaving the galaxy. Each time a photon goes through one of these events, it changes in direction and frequency because of the velocity distribution of the hydrogen atoms. After escaping, the ${\rm Ly}\alpha$ line profile (frequency distribution) is greatly influenced by the gas geometry, density, and bulk velocity.
The fact that ${\rm Ly}\alpha$ observables probe gas from the ISM to the IGM holds tremendous potential. As such, ${\rm Ly}\alpha$ spectra are frequently used to probe the porosity of the ISM and as a proxy for ionizing photon escape (Izotov et al., 2016; Dijkstra et al., 2016; Verhamme et al., 2017, e.g.,). Furthermore, the existence of a dominant red peak in the ${\rm Ly}\alpha$ spectra (e.g. Erb et al., 2014; Trainor et al., 2015) is interpreted as a widespread presence of galactic outflows.
On the scale of the CGM, ${\rm Ly}\alpha$ reveals the gas properties through the shape of the surface brightness profiles on scales of tens of kpc for individual (e.g., Wisotzki et al., 2016) and even hundreds of kpc for stacked observations (Steidel et al., 2011; Wisotzki et al., 2018; Lujan Niemeyer et al., 2022). In addition, thanks to the advent of integral field spectrographs such as MUSE (Bacon et al., 2010) or KCWI (Martin et al., 2010), one can now also study of the emergent ${\rm Ly}\alpha$ spectrum changes as a function of projected position (e.g. Erb et al., 2018; Leclercq et al., 2017) revealing, for instance, the transition from galactic out- to intergalactic inflows (Erb et al., 2018; Li et al., 2022; Guo et al., 2024).
On even larger scales, the ${\rm Ly}\alpha$ photons emerging from the galaxy are shaped by the IGM, which contains neutral hydrogen in (mostly self-shielding) lumps. Thus, further radiative transfer processes influence the ${\rm Ly}\alpha$ line profile. However, because photons interacting with the IGM gas change direction and are scattered out of the line of sight – and the probability of scattering (back) into the line of sight is negligible – one can treat these large-scale effects as effective absorption. Importantly, in this process, photons are redshifted progressively by the Hubble flow as they travel through the IGM. This causes the IGM to attenuate higher frequencies of the galactic spectrum as photons advance (e.g. Laursen et al., 2011a).
@@
While this ability of ${\rm Ly}\alpha$ to probe neutral hydrogen on the ISM, CGM, and IGM scales provides the community with a potential treasure trove of information, this highly multiscale process naturally comes with many degeneracies. For instance, does the absence of a blue peak imply an ISM outflow or a neutral IGM? To unravel this degeneracy, we implemented the ability to reconstruct the transmission and the ‘galactic’ ${\rm Ly}\alpha$ spectra (shaped by the ISM and CGM) into the Python package zELDA  [^1] Github: https://github.com/sidgl/zELDA_II. Tutorials and documentation on zELDA are also available at https://zelda-ii.readthedocs.io/index.html. (redshift Estimator for Line profiles of Distant Lyman-Alpha emitters). To fit observed ${\rm Ly}\alpha$ line profiles, zELDA is now able to disentangle between the galactic and the IGM components of a ${\rm Ly}\alpha$ line profile through artificial neural network models trained with mock ${\rm Ly}\alpha$ spectra (see, Gurung-Lopez et al. (in prep.) for a detailed description and testing of the algorithm or Sect. 2 for a summary).
How the ISM, CGM, and IGM affect the observed ${\rm Ly}\alpha$ spectra and the ability to disentangle the ‘galactic’ from the intergalactic effects are visualized in Fig. 1. Here, we show in the top panels how the ${\rm Ly}\alpha$ line is affected by absorption in the IGM (as described above) and in the bottom panels how zELDA manages to reconstruct the ‘galactic’ line shape relatively well.
@@
This work is structured as follows: in Sect. 2 we briefly introduce zELDA’s pipeline for recovering the galactic component from IGM attenuated ${\rm Ly}\alpha$ line profiles. We introduce the observational data set used in this work in Sect. 3. Next, we apply zELDA’s methodology to the observational data set in Sect. 4. Finally, we draw our conclusions in Sect. 5.
@@
In this work, we draw the intergalactic transmission curves and ${\rm Ly}\alpha$ line profiles in the rest frame difference to the ${\rm Ly}\alpha$ wavelength ($\Delta\lambda_{0}$). Redshift accuracy is provided in the same units. Typically, this quantity is expressed in velocity units, instead, as $\Delta v=c\Delta\lambda_{0}/\lambda_{\rm Ly\alpha}\approx(247\,{\mathrm{km}\,\mathrm{s}^{-1}})\times\Delta\lambda/1$Å, where $c$ is the void speed of light and $\lambda_{\rm Ly\alpha}\approx 1215.67$Å.
@@
Furthermore, we refer to the ${\rm Ly}\alpha$ IGM escape fraction of a source as the ratio between the flux emerging from the ISM/CGM and that observed after the IGM along the observer’s line of sight. This definition can be understood as the transmission of IGM around ${\rm Ly}\alpha$ and through the line of sight. In general, in the IGM, ${\rm Ly}\alpha$ photons are not destroyed by dust grains, although they are scattered out of the line of sight. This effectively causes a reduction of the observed ${\rm Ly}\alpha$ flux.
@@ 2 Reconstructing attenuated Lyman-$\alpha$ emission lines
In this section, we detail our methodology to reconstruct observed ${\rm Ly}\alpha$ line profiles attenuated by the intergalactic medium and, simultaneously, estimate their ${\rm Ly}\alpha$ IGM escape fraction. Our methodology is based on the Python package zELDA. Here, we briefly introduce zELDA’s pipeline, which is fully detailed in ZP24.
@@
To fit the observed ${\rm Ly}\alpha$ line profiles, we use the open-source Python package zELDA (Gurung-López et al., 2019a, 2021, 2022). zELDA’s thin shell model is integrated by linear interpolation in a precomputed grid of line profiles provided by the full Monte Carlo radiative transfer code LyaRT (Orsi et al., 2012). The grid contains 3,132,000 nodes in the five dimensional space of bulk expansion velocity $V_{\rm exp}$ $\in[0,1000]kms^{-1}$, neutral hydrogen column density $N_{\rm H}$ $\in[10^{17},10^{21.5}]cm^{-2}$, dust optical depth $\tau_{a}$ $\in[0.0001,0.0]$, the intrinsic equivalent width $EW_{\rm in}$ $\in[0.1,1000]\AA{}$ and line width $W_{\rm in}$ $\in[0.01,6]\AA{}$ of the ${\rm Ly}\alpha$ emission before entering the thin shell. Further information on the grid and its accuracy can be found in Gurung-López et al. (2022).
@@
The IGM transmission curves used to train the IGM+z and IGM-z models are described in detail in Byrohl & Gronke (2020); Byrohl et al. (2021). These are computed in the IllustrisTNG100 simulation (Naiman et al., 2018; Nelson et al., 2019; Marinacci et al., 2018; Pillepich et al., 2018; Springel et al., 2018). The ${\rm Ly}\alpha$ radiative transfer was computed using an updated version of the code ILTIS (Behrens et al., 2019; Byrohl & Gronke, 2020; Byrohl et al., 2021). An IGM transmission curve is obtained for every halo heavier than $5\times 10^{9}M_{\odot}$ in 1000 different lines of sight. This analysis was performed in six redshift snapshots at $z=$0.0, 1.0, 2.0, 3.0, 4.0 and 5.0.
@@
zELDA includes three models based on artificial neural networks. For all of them, the input contains an observed ${\rm Ly}\alpha$ line profile. The models IGM+z and IGM-z take into account the IGM attenuation and provide an unattenuated best-fitting shell model. IGM+z is conceived as a model with the IGM redshift evolution of IllustrisTNG100. Meanwhile, IGM-z is built to be redshift independent. The third model, NoIGM, does not consider the IGM and fits a shell model directly to the observed line profile.
@@
  • IGM-z: The training set also includes mock line profiles attenuated by the IGM. Although, in contrast to IGM+z, we use the re-calibrated  [^2] As shown in ZP24, Byrohl & Gronke (2020) IGM transmission curves were fine-tuned so that they match the mean evolution of the redshift of the IGM transmission by Faucher-Giguère et al. (2008). Thus, the IGM transmission curves used in IGM+z and IGM-z are slightly different. transmission curves. These are drawn randomly without considering the mock source redshift. Thus, there is no evolution of the IGM through redshift in this training set. There will be ${\rm Ly}\alpha$ basically unattenuated (typical of $z=0$) and greatly attenuated (typical of high $z$) at all redshifts. Finally, unlike IGM+z, IGM-z does not include the proxy redshift in the input.
The output of the three models is basically the same: the shell model parameters with their uncertainty, the redshift of the source, and the IGM escape fraction 2Å around the ${\rm Ly}\alpha$ wavelength, $f_{\rm esc}^{4\AA}$ (in the case of IGM+z and IGM-z).
The motivation for using IGM+z and IGM-z is to verify the validity of our results. While IGM+z is trained with the evolution of IGM transmission with redshift of IllustrisTNG100, IGM-z lacks it. Comparison between these two models can highlight possible biases in our methodology in redshift-dependent trends, such as the evolution of the IGM escape fraction. In principle, if analyzing IGM+z predictions, we find a redshift dependence trend missing from IGM-z data, then it could be induced by our training set rather than being real. Meanwhile, we show the results of NoIGM, a model without IGM attenuation for comparison.
Figure 2: Nine examples of zELDA’s prediction on line profiles displayed in the rest frame. The observed line profile is shown in dark grey with its $1\sigma$ uncertainty in light grey. zELDA’s reconstruction using the models IGM+z, IGM-z and NoIGM are displayed in blue, green and yellow, respectively. The redshift of the source is shown in the top right corner. zELDA’s estimations of $f_{\rm esc}^{4\AA}$ given by the IGM+z and IGM-z models are shown in blue and green, respectively. The central grey-shaded region shows the wavelength interval of $f_{\rm esc}^{4\AA}$. The estimated outflow parameters are indicated in the table with each subplot. $V_{\rm exp}$ is given in  km s-1, $N_{\rm H}$ in $cm^{-2}$ and $EW_{\rm in}$ and $W_{\rm in}$ in Å. All the fitted line profiles are also shown in Appendix C.
3 Description of observational data
The observational ${\rm Ly}\alpha$ line profiles used in this work are stored in the Lyman alpha Spectral Database (LASD, Runnholm et al., 2021) [^3] http://lasd.lyman-alpha.com, which currently contains 349 ${\rm Ly}\alpha$ emission lines between redshift 0 and 6.0. We remove from our analysis
  1. 1.
    ${\rm Ly}\alpha$ line profiles with a very steep continuum around ${\rm Ly}\alpha$,
@@ 3 Description of observational data
    spectra that apparently contain the ${\rm Ly}\alpha$ emission line of several sources.
This leaves a total of 313 ${\rm Ly}\alpha$ line profiles from redshift 0 to 6.0. All the spectra are displayed in Appendix C.
@@
The ${\rm Ly}\alpha$ line profiles from LASD were obtained mainly from two experiments: i) the Cosmic Origins Spectrograph (COS Green et al., 2012) on board the Hubble Space Telescope (HST) and ii) the MUSE-WIDE survey Urrutia et al. (2019); Herenz et al. (2017). ${\rm Ly}\alpha$ line profiles obtained via HST were observed in the General Observers (GO): GO 11522 and 12027 (PI: Green, Salzer et al., 2001; Wofford et al., 2013), GO11727 and 13017 (PI: Heckman, Heckman et al., 2011, 2015), GO 12269 (PI: Scarlata, Songaila et al., 2018), GO 12583 (PI: Hayes Hayes et al., 2014; Rivera-Thorsen et al., 2015), GO12928 (PI: Henry Henry et al., 2015), GO 13293 and 14080 (PI: Jaskot Jaskot & Oey, 2014; Jaskot et al., 2017), GO 14201 (PI: Malhotra Yang et al., 2017) and GO 13744 (PI: Thuan Izotov et al., 2016, 2018, 2020).
@@
HST-COS and MUSE-WIDE exhibit very different spectral quality. On one side, the spectra obtained by HST have excellent spectral quality ($R\sim 16,000$). In particular, the ${\rm Ly}\alpha$ line profiles exhibit excellent wavelength sampling, as $\Delta\lambda_{\rm Pix}$ might take two particular values: 0.0598Å and 0.0735Å. This is given by the two medium resolution gratings G130M and G160M used to observe the sources. The spectral resolution is also high quality. $W_{\rm g}$ and ranges from 0.073Å to 0.10Å with median 0.085Å. The signal-to-noise ratio of the maximum of the line, $S/N_{p}$, spawns a wide range from $S/N_{p}$ =6.5 to $\sim 400$ with a median of $\sim$38 and only a $\sim 11\%$ of sample exhibit $S/N_{p}$ $<15.0$. On the other hand, considering MUSE data, the spectral quality is more limited ($R\sim 3000$), $\Delta\lambda_{\rm Pix}$ =1.25Å. Meanwhile, $W_{\rm g}$ changes from 1.25Å at $z=3$ to 4.0Å at $z=6.0$. We only consider ${\rm Ly}\alpha$ line profiles such as $S/N_{p}$5.0. The maximum $S/N_{p}$ is 35.0, and the median is 7.8.
In ZP24 several spectral quality configurations are explored. Most of the line profiles obtained by HST have a similar quality to the best-case scenario studied in the previous sections, and these ${\rm Ly}\alpha$ line profiles are presented in sections 4.2 and 4.3 in ZP24. Meanwhile, MUSE spectral quality is similar to the middle and worst cases in the same sections of ZP24. It is shown in ZP24 that for MUSE and HST spectral quality, zELDA can reconstruct the stacked ${\rm Ly}\alpha$ line profile, and in particular if there is evolution, as well as the $f_{\rm esc}^{4\AA}$ redshift evolution.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Figure 1: Illustration of the impact of the IGM attenuation in the galactic ${\rm Ly}\alpha$ line profile. Through all the panels, we show the intrinsic ${\rm Ly}\alpha$ line profile emerging from the galaxy in thin black and the IGM transmission curve in yellow. In the top panels we show the galactic line profile obscured intrinsically (top left, purple), the same profile after traveling a short distance in the IGM (blue, center left) as well as further through the IGM (green, top right), and its shape when it reaches the observer (red). The bottom panel show the zELDA reconstruction (purple) using as input the observed line profile (red), along with the line structure terminology used in this work.
@@
The emergent ${\rm Ly}\alpha$ line profile results from complex radiative-transfer processes in the interstellar, circumgalactic, and intergalactic media (ISM, CGM, and IGM; see Dijkstra 2017 for a review). Because of the large scattering cross-section of neutral hydrogen, ${\rm Ly}\alpha$ photons are absorbed and reemitted many times before escaping the galaxy. Each scattering event changes both the direction and the frequency of the photon according to the velocity distribution of the gas. Consequently, the observed ${\rm Ly}\alpha$ profile is strongly shaped by the geometry, density, and bulk motions of the gas.
The fact that ${\rm Ly}\alpha$ observables probe gas from the ISM to the IGM holds tremendous potential. As such, ${\rm Ly}\alpha$ spectra are frequently used to investigate the porosity of the ISM and as a proxy for ionizing photon escape (e.g., Izotov et al., 2016; Dijkstra et al., 2016; Verhamme et al., 2017). Furthermore, the existence of a dominant red peak in the ${\rm Ly}\alpha$ spectra (e.g. Erb et al., 2014; Trainor et al., 2015) is interpreted as a widespread presence of galactic outflows.
On the scale of the CGM, ${\rm Ly}\alpha$ reveals the gas properties through the shape of the surface brightness profiles on scales of tens of kpc for individual (e.g., Wisotzki et al., 2016) and even hundreds of kpc for stacked observations (Steidel et al., 2011; Wisotzki et al., 2018; Lujan Niemeyer et al., 2022). In addition, thanks to the advent of integral field spectrographs such as MUSE (Bacon et al., 2010) or KCWI (Martin et al., 2010), one can now also study the emergent ${\rm Ly}\alpha$ spectrum changes as a function of projected position (e.g. Erb et al., 2018; Leclercq et al., 2017), revealing, for instance, the transition from galactic out- to intergalactic inflows (Erb et al., 2018; Li et al., 2022; Guo et al., 2024).
At even larger scales, the ${\rm Ly}\alpha$ photons emerging from the galaxy are shaped by the IGM, which contains neutral hydrogen in (mostly self-shielding) lumps. Thus, further radiative transfer processes influence the ${\rm Ly}\alpha$ line profile. However, because photons interacting with the IGM gas change direction and are scattered out of the line of sight – and the probability of scattering (back) into the line of sight is negligible – one can treat these large-scale effects as effective absorption. Importantly, in this process, photons are redshifted progressively by the Hubble flow as they travel through the IGM. This redshifting causes the IGM to attenuate higher frequencies of the galactic spectrum as photons advance (e.g. Laursen et al., 2011a).
@@
While this ability of ${\rm Ly}\alpha$ to probe neutral hydrogen on the ISM, CGM, and IGM scales provides the community with a potential treasure trove of information, this highly multiscale process naturally comes with many degeneracies. For instance, does the absence of a blue peak imply an ISM outflow or a neutral IGM? To unravel this degeneracy, we implemented the ability to reconstruct the transmission and the ‘galactic’ ${\rm Ly}\alpha$ spectra (shaped by the ISM and CGM) into the Python package zELDA  [^1] Github: https://github.com/sidgl/zELDA_II. Tutorials and documentation on zELDA are also available at https://zelda-ii.readthedocs.io/index.html. (redshift Estimator for Line profiles of Distant Lyman-Alpha emitters). To fit observed ${\rm Ly}\alpha$ line profiles, zELDA is now able to disentangle between the galactic and the IGM components of a ${\rm Ly}\alpha$ line profile through artificial neural network models trained with mock ${\rm Ly}\alpha$ spectra (see Gurung-Lopez et al. (2025) for a detailed description and testing of the algorithm or Sect. 2 for a summary).
This work aims to demonstrate on real data, i.e., actually observed spectra, how one can disentangle between the IGM and the ISM/CGM effects shaping the ${\rm Ly}\alpha$ spectral shapes. With this separation of effects at hand, we will also study the evolving impact of the IGM on spectral shapes, or vice versa study the IGM evolution purely using ${\rm Ly}\alpha$ spectra. A similar exercise has been done in previous work using stacked line profiles (Hayes et al., 2011). The novelty of this work is that this will be the first time doing so using individual spectra and estimating the ${\rm Ly}\alpha$ escape fraction in the separate sources. While the current work represents only the first step on utilizing this potential treasure trove of information (and thus, is to some degree a proof-of-concept study), other applications such as spatial tomography of the IGM are thinkable.
@@
This work is structured as follows: in Sect. 2 we briefly introduce zELDA’s pipeline for recovering the galactic component from IGM attenuated ${\rm Ly}\alpha$ line profiles. We introduce the observational data set used in this work in Sect. 3. Next, we apply zELDA’s methodology to the observational data set in Sect. 4 and measure the mean IGM ${\rm Ly}\alpha$ escape fraction as a function of redshift. In Sect. 5 we validate our results by comparing them to the literature. Finally, we draw our conclusions in Sect. 6.
@@
In this work, we draw the intergalactic transmission curves and ${\rm Ly}\alpha$ line profiles in the rest frame difference to the ${\rm Ly}\alpha$ wavelength ($\Delta\lambda_{0}$). Redshift accuracy is provided in the same units. Typically, this quantity is expressed in velocity units, instead, as $\Delta v=c\Delta\lambda_{0}/\lambda_{\rm Ly\alpha}\approx(247\,{\mathrm{km}\,\mathrm{s}^{-1}})\times\Delta\lambda/1$Å, where $c$ is the speed of light in vacuum and $\lambda_{\rm Ly\alpha}\approx 1215.67$Å.
@@
Furthermore, we refer to the ${\rm Ly}\alpha$ IGM escape fraction of a source as the ratio between the flux emerging from the ISM/CGM and that observed after the IGM along the observer’s line of sight. This definition can be understood as the transmission of IGM around ${\rm Ly}\alpha$ and through the line of sight, convoluted with the line profile shape. In general, in the IGM, ${\rm Ly}\alpha$ photons are not destroyed by dust grains, although they are scattered out of the line of sight. This effectively causes a reduction of the observed ${\rm Ly}\alpha$ flux.
@@ 2 Reconstructing attenuated Lyman-$\alpha$ emission lines
In this section, we describe our approach to reconstructing observed ${\rm Ly}\alpha$ line profiles attenuated by the intergalactic medium (IGM) and to estimating their ${\rm Ly}\alpha$ IGM escape fraction. Our methodology is based on the Python package zELDA. Here, we briefly summarize zELDA’s pipeline, which is fully detailed in Gurung-Lopez et al. (2025), hereafter Z25.
@@
To fit the observed ${\rm Ly}\alpha$ line profiles, we use the open-source Python package zELDA (Gurung-López et al., 2019a, 2021, 2022). zELDA ’s thin-shell model is implemented via linear interpolation within a precomputed grid of line profiles generated by the full Monte Carlo radiative transfer code LyaRT (Orsi et al., 2012). The grid contains 3,132,000 nodes spanning a five-dimensional parameter space: bulk expansion velocity $V_{\rm exp}$ $\in[0,1000]\,\mathrm{km\,s^{-1}}$, neutral hydrogen column density $N_{\rm H}$ $\in[10^{17},10^{21.5}]\,\mathrm{cm^{-2}}$, dust optical depth $\tau_{a}$ $\in[0.0001,1.0]$, intrinsic equivalent width $EW_{\rm in}$ $\in[0.1,1000]\,\AA$ , and line width $W_{\rm in}$ $\in[0.01,6]\,\AA$ of the ${\rm Ly}\alpha$ emission before entering the thin shell. Further details on the grid and its accuracy can be found in Gurung-López et al. (2022).
@@
The IGM transmission curves used to train the IGM+z and IGM-z models are described in detail in Byrohl & Gronke (2020); Byrohl et al. (2021). These were computed within the TNG100 simulation (Naiman et al., 2018; Nelson et al., 2019; Marinacci et al., 2018; Pillepich et al., 2018; Springel et al., 2018). The ${\rm Ly}\alpha$ radiative transfer was calculated using an updated version of the ILTIS code (Behrens et al., 2019; Byrohl & Gronke, 2020; Byrohl et al., 2021). For each halo more massive than $5\times 10^{9}M_{\odot}$, transmission curves were obtained along 1000 random sightlines, at six redshift snapshots (z=0.0, 1.0, 2.0, 3.0, 4.0, 5.0z=0.0,\,1.0,\,2.0,\,3.0,\,4.0,\,5.0).
@@
zELDA incorporates three models based on artificial neural networks (ANNs). Each model takes as input an observed ${\rm Ly}\alpha$ line profile. Both IGM+z and IGM-z account for IGM attenuation and aim to recover the unattenuated best-fitting shell model. IGM+z is trained to capture the redshift evolution of the IGM according to TNG100, while IGM-z is constructed to be redshift-independent. These two models therefore complement each other when measuring redshift-dependent properties. The third model, NoIGM, completely ignores IGM effects and fits a shell model directly to the observed line profile.
The three artificial neural network models are trained with mocked ${\rm Ly}\alpha$ line profiles produced with zELDA and the IGM transmission curves from Byrohl & Gronke (2020); Byrohl et al. (2021). The mock line profiles cover homogeneously the outflow parameters of the shell model. The training sets cover a wide range of spectral quality, from resolution of $W_{\rm g}$ =0.1Å to $W_{\rm g}$ =4.0Å and signal to noise of the peak of the line between $S/N_{p}$ =5.0 and $S/N_{p}$ =20.0. IGM+z, IGM-z, and NoIGM use a different training set and input. In summary:
  • IGM-z: The training set also includes mock line profiles attenuated by the IGM. Although, in contrast to IGM+z, we use the re-calibrated  [^2] As shown in Z25, Byrohl & Gronke (2020) IGM transmission curves were fine-tuned so that they match the mean evolution of the redshift of the IGM transmission ($\langle T_{\rm IGM}\rangle$) by Faucher-Giguère et al. (2008). Thus, the IGM transmission curves used in IGM+z and IGM-z are slightly different. transmission curves. These are drawn randomly without considering the mock source redshift. Thus, there is no evolution of the IGM through redshift in this training set. There will be ${\rm Ly}\alpha$ basically unattenuated (typical of $z=0$) and greatly attenuated (typical of high $z$) at all redshifts. Finally, unlike IGM+z, IGM-z does not include the proxy redshift in the input.
All three models provide the same type of output: the shell-model parameters with uncertainties, the estimated redshift, and the ${\rm Ly}\alpha$ IGM escape fraction within $\pm 2\,\AA$ of the line center, $f_{\rm esc}^{4\AA }$ (for IGM+z and IGM-z). By definition, $f_{\rm esc}^{4\AA }$ is obtained as the ratio of the IGM-convolved ${\rm Ly}\alpha$ flux to the intrinsic flux. Since the IGM at $z<4$ primarily affects the blue side of ${\rm Ly}\alpha$, $f_{\rm esc}^{4\AA }$ specifically measures absorption in the $\pm 2\,\AA$ window, and thus, while related, differs from the mean IGM transmission.
The use of both IGM+z and IGM-z enables us to test the robustness of our measurements. IGM+z encodes the expected redshift evolution of IGM transmission, whereas IGM-z does not. Comparing them allows us to identify potential biases in redshift-dependent trends, such as spurious variations in the IGM escape fraction induced by the training set. If IGM+z predicts a redshift trend absent in IGM-z, the effect is likely artificial and it could be attributed to our training set rather than being authentic. For reference, we also include results from NoIGM, which ignores IGM attenuation entirely.
This paper aims to evaluate how the ${\rm Ly}\alpha$ IGM escape fraction evolve with redshift. While the mean IGM transmission is well known to evolve, the redshift dependence of the ${\rm Ly}\alpha$ escape fraction has not been directly measured before. On the other hand, $\langle f_{\rm esc}^{4\AA }\rangle$ depends on the local IGM environment (typically denser and more neutral than the field) as well as on the ${\rm Ly}\alpha$ line profile shape. Thus, it is unclear whether the IGM+z training set reproduces its evolution accurately. We therefore include the IGM-z model as a redshift-independent control.
When the spectral quality is sufficiently high, both IGM-z and IGM+z recover $f_{\rm esc}^{4\AA }$ and the outflow parameters reliably, as shown in Z25. At low quality, however, the models behave differently due to their training sets. In particular, as quality decreases, the ANNs tend to return values that reflect prior distributions rather than true inference. Since IGM+z is forced to encode redshift evolution, it may reproduce trends even when absent in the data, while IGM-z will not. The close agreement between their predictions in the results section is reassuring, as it indicates that the information content of the observed spectra is sufficient for accurate inference rather than being dominated by training priors.
3 Description of observational data
The observational ${\rm Ly}\alpha$ line profiles used in this work are stored in the Lyman Alpha Spectral Database (LASD, Runnholm et al., 2021) [^3] http://lasd.lyman-alpha.com, which currently contains 349 ${\rm Ly}\alpha$ emission lines in the redshift range $0<z<6.0$. We exclude from our analysis:
  1. 1.
@@ 3 Description of observational data
    spectra that apparently contain ${\rm Ly}\alpha$ emission from multiple sources.
After these cuts, our final sample comprises 313 ${\rm Ly}\alpha$ line profiles spanning $0<z<6.0$. All spectra are shown in Appendix F.
@@
The ${\rm Ly}\alpha$ line profiles from LASD were obtained mainly from two experiments: i) the Cosmic Origins Spectrograph (COS Green et al., 2012) on board the Hubble Space Telescope (HST) and ii) the MUSE-WIDE survey (Urrutia et al., 2019; Herenz et al., 2017). ${\rm Ly}\alpha$ line profiles obtained via HST were observed in the General Observers (GO): GO 11522 and 12027 (PI: Green, Salzer et al., 2001; Wofford et al., 2013), GO11727 and 13017 (PI: Heckman, Heckman et al., 2011, 2015), GO 12269 (PI: Scarlata, Songaila et al., 2018), GO 12583 (PI: Hayes Hayes et al., 2014; Rivera-Thorsen et al., 2015), GO12928 (PI: Henry Henry et al., 2015), GO 13293 and 14080 (PI: Jaskot Jaskot & Oey, 2014; Jaskot et al., 2017), GO 14201 (PI: Malhotra Yang et al., 2017) and GO 13744 (PI: Thuan Izotov et al., 2016, 2018, 2020).
@@
HST-COS and MUSE-WIDE exhibit very different spectral quality. The spectra obtained by HST have excellent spectral quality ($R\sim 16,000$). In particular, the ${\rm Ly}\alpha$ line profiles exhibit excellent wavelength sampling, as $\Delta\lambda_{\rm Pix}$ might take two particular values: 0.0598Å and 0.0735Å, depending on the medium-resolution gratings G130M and G160M. The spectral resolution is also of high quality. $W_{\rm g}$ and ranges from 0.073Å to 0.10Å with median 0.085Å. The signal-to-noise ratio of the maximum of the line, $S/N_{p}$, spans a wide range from $S/N_{p}$ =6.5 to $\sim 400$ with a median of $\sim$38 and only a $\sim 11\%$ of the sample exhibit $S/N_{p}$ <15.0<15.0. Considering MUSE data, the spectral quality is more limited ($R\sim 3000$), $\Delta\lambda_{\rm Pix}$ =1.25Å. Meanwhile, $W_{\rm g}$ changes from 1.25Å at $z=3$ to 4.0Å at $z=6.0$. We only consider ${\rm Ly}\alpha$ line profiles such as $S/N_{p}$ >>5.0. The maximum $S/N_{p}$ is 35.0, and the median is 7.8.
Figure 2: Nine examples of zELDA’s prediction on line profiles displayed in the rest frame. The observed line profile is shown in dark grey with its $1\sigma$ uncertainty in light grey. zELDA’s reconstruction using the models IGM+z, IGM-z, and NoIGM is displayed in blue, green, and yellow, respectively. The redshift of the source is shown in the top right corner. zELDA’s estimations of $f_{\rm esc}^{4\AA }$ given by the IGM+z and IGM-z models are shown in blue and green, respectively. The central grey-shaded region shows the wavelength interval of $f_{\rm esc}^{4\AA }$. The estimated outflow parameters are indicated in the table with each subplot. $V_{\rm exp}$ is given in  km s-1, $N_{\rm H}$ in $cm^{-2}$ and $EW_{\rm in}$ and $W_{\rm in}$ in Å. All the fitted line profiles are also shown in Appendix F.
In Z25, different spectral-quality configurations were explored. Most HST line profiles correspond to the best-case scenarios described there, while the MUSE spectra match intermediate and worst-case configurations. It was demonstrated in Z25 that zELDA can successfully reconstruct stacked ${\rm Ly}\alpha$ line profiles for both MUSE and HST qualities, and can reliably track the redshift evolution of $f_{\rm esc}^{4\AA }$.
The $S/N_{p}$ thresholds applied here were chosen to ensure accurate recovery of the stacked line profile, $f_{\rm esc}^{4\AA }$, and its mean $\langle f_{\rm esc}^{4\AA }\rangle$ across redshift, as shown in Z25. Specifically, for HST sources zELDA recovers the stacked line profile with a Kolmogorov–Smirnov (KS) estimator better than 0.04.444The KS estimator is computed between the reconstructed and true intrinsic ${\rm Ly}\alpha$ line profiles for 10,000 mock spectra spanning the full range of zELDA outflow parameters, IGM absorption levels, and spectral-quality configurations. For MUSE sources the stacked line profile is recovered with KS <0.11<0.11. As shown in Z25 (Sect. 4.3), with the spectral qualities of both MUSE and HST the redshift evolution (or lack thereof) of the stacked ${\rm Ly}\alpha$ profile is recovered robustly. Regarding $\langle f_{\rm esc}^{4\AA }\rangle$, for HST data both IGM+z and IGM-z retrieve unbiased values within 3%, whereas for MUSE-quality spectra $\langle f_{\rm esc}^{4\AA }\rangle$ may be biased by up to 10%. [^5] The precision of $f_{\rm esc}^{4\AA }$ estimates is evaluated using mock ${\rm Ly}\alpha$ profiles generated with zELDA under varying spectral qualities and IGM absorption levels.
[…]